### v1

@@ 4.3.4 Putting it all together
Case 2: $\underline{X}\leq\frac{B}{N}\leq\overline{X}$. We have: $|X_{t,\theta}^{alg}-B/N|=\max\mathopen{}\mathclose{{}\left\{\frac{B}{N}-\underline{X},\overline{X}-\frac{B}{N}}\right\}\leq\overline{X}-\underline{X}=L_{T}.$
@@
where the inequality follows from $X_{t_{0}}^{alg}=\underline{X}$, and the threshold condition for allocating $\overline{X}$.
@@
$\leq 2\beta_{avg}\sqrt{2\tilde{c}|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}\min\mathopen{}\mathclose{{}\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}}\right\},$ (14)
@@
Finally, consider $\overline{P}_{t_{0}}+\textsc{PUA}_{<t_{0}}^{alg}$. By construction, $\overline{P}_{t_{0}}\leq\overline{P}_{1}=\overline{\Delta}(\underline{X})$. To upper bound $\textsc{PUA}_{<t_{0}}^{alg}$, we consider the process that allocates $\underline{X}$ in each period to all arrivals. Let $B_{t}(\underline{X})$ denote the quantity of remaining items under this process, and $\mathcal{B}_{t}(\underline{X})$ the set of remaining items. We use $\textsc{PUA}_{t}(\underline{X})$ to denote the quantity of unallocated items that perish at the end of period $t$ under this process, and $\textsc{PUA}_{<t}(\underline{X})$ those that perished before the end of period $t$. The following lemma allows us to tractably bound $\textsc{PUA}_{<t_{0}}^{alg}$ via this process. We defer its proof to Appendix B.5.4.
Lemma 4.13.
@@
where the second inequality follows from $\overline{\Delta}(\underline{X})\leq B-\overline{N}\underline{X}=B-\overline{N}\mathopen{}\mathclose{{}\left(\frac{B}{\overline{X}}-\mathcal{L}^{\textsf{perish}}}\right)=\overline{N}\mathcal{L}^{\textsf{perish}}$, and the last inequality uses the definition of $\overline{N}$. Putting bounds (4.3.4), (15) and (16) together, we obtain:
$\Delta_{\text{\it efficiency}}$ $\leq 2\beta_{avg}\sqrt{2\tilde{c}|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}\min\mathopen{}\mathclose{{}\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}}\right\}+L_{T}|\Theta|(\mu_{max}+\rho_{\max})$
@@
Using the fact that $L_{T}=o(1)$, we obtain the final bound on efficiency. ∎
@@ 5 Numerical experiments
In this section we study the practical performance of Perishing-Guardrail via an extensive set of numerical experiments. We first consider one of the most popular (and aggressive) models of perishing: geometrically distributed perishing times. For this tractable perishing process, we establish distribution-dependent bounds on the $\sigma$-induced loss $\mathcal{L}^{\textsf{perish}}$, and empirically explore the dependence of the envy-efficiency trade-off on the perishing rate. We leverage these empirical trade-off curves to provide guidance on how to select the envy parameter $L_{T}$, and compare the performance of Perishing-Guardrail to its perishing-agnostic counterpart (Sinclair et al., , 2022). We moreover demonstrate the robustness of our algorithm on a real-world dataset on ginger perishability (Keskin et al., , 2022). We conclude our numerical study by considering the non-i.i.d. perishing setting to gain insights into the choice of allocation schedule. Code available at https://github.com/seanrsinclair/Online-Resource-Allocation
5.1 Geometric perishing
@@
Consider the setting in which each available unit perishes independently with probability $p$ in each period, i.e., $T_{b}\sim\text{Geometric}(p)$, for all $b\in\mathcal{B}$. Throughout the section, we assume $|\Theta|=1$, as our insights are invariant to the number of types. Since perishing times are identically distributed, the allocation order $\sigma$ does not have any impact on the performance of the algorithm; hence, in the remainder of this section we assume $\sigma$ is the identity ordering.
5.1.1 Quantifying the unavoidable perishing loss
We first investigate the impact of perishability by characterizing the lower guardrail $\underline{X}$ as a function of the perishing rate $p.$ As in Section 4.2, we instantiate our theoretical bounds assuming $B=T,N_{t}=1$ for all $t\in[T]$. In this case, Proposition 4.6 implies that $p\leq\frac{1}{T}$ is necessary to guarantee $\delta$-offset-expiry, for nontrivial values of $\delta$. Proposition 5.1 below provides a lower bound on $\underline{X}$ for this setting. We defer its proof to Section B.6.
Proposition 5.1.
Suppose $T_{b}\sim\emph{Geometric}(p)$ for all $b\in\mathcal{B}$, with $p\leq 1/T$. Then, the perishing process is $\delta$-offset-expiring for any $\delta\geq 2\log T\cdot\frac{Tp}{(1-Tp)^{2}}$. Moreover, $\underline{X}\geq 1-3Tp-\frac{\log(3\log(T)/\delta)}{T}$ for any ordering $\sigma$.
Figure 3: Maximum feasible allocation $\underline{X}$ vs. $\alpha$, for $T_{b}\sim\emph{Geometric}(T^{-(1+\alpha)})$, $B=200$, $T=150$, $N_{t}$ drawn from a truncated normal distribution $\mathcal{N}(2,0.25)$, and $\delta=1/T$. Here, $\underline{X}$ was calculated via line search, with Monte Carlo simulation used to estimate $\overline{\Delta}(X)$ for each value of $X$. The dashed line represents the “naive” allocation $B/\overline{N}=0.89$ which ignores possible perishing, and the green line is the curve of best fit to $\underline{X}$.
Proposition 5.1 establishes that, in the worst case, $\underline{X}$ decays linearly in the rate $p$ at which goods spoil. This highlights the extent to which the commonly used (and practically pessimistic) geometric model limits both the kinds of perishable goods selected before allocation, as well as the rate at which a decision-maker can allocate goods. Letting $p=T^{-(1+\alpha)}$, $\alpha\in(0,1)$, Proposition 5.1 implies that $\mathcal{L}^{\textsf{perish}}$ is on the order of $T^{-\alpha}$. Alternatively, if a decision-maker wants no more than $T^{-\alpha}$ loss relative to the proportional allocation, Proposition 5.1 provides an upper bound of $T^{-(1+\alpha)}$ on the (exogenous) rate at which goods spoil.
We validate the scaling of $\underline{X}$ numerically in Fig. 3, for an instance where the decision-maker also faces demand uncertainty. We observe that $\underline{X}$ is concave increasing in $\alpha$, and that our lower bound on $\underline{X}$ in the idealized setting provides a good fit for $\underline{X}$, even under demand uncertainty. For $\alpha$ close to 0 (i.e., $p\sim 1/T$), $\underline{X}\approx 0.4$, less than half of the “naive” no-perishing allocation, $B/\overline{N}$. For $\alpha=1$, $\underline{X}\approx B/\overline{N}$. Note that, for $\alpha>1$, $\underline{X}$ is limited by the confidence bound $\log(3\log T/\delta)/T$ in Proposition 5.1. Plugging the lower bound on $\delta$ into this term, this implies that, even under no demand uncertainty, $\underline{X}$ incurs a loss on the order of $1/T$ relative to the proportional allocation.
@@ 5.1.1 Quantifying the unavoidable perishing loss
Figure 4: Empirical trade-off between $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right]$ and $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$. The points on the trade-off curve correspond to increasing values of $L_{T}$, from left to right. Static-$\frac{B}{\overline{N}}$ and Static-$\underline{X}$ respectively correspond to Vanilla-Guardrail and Perishing-Guardrail for $L_{T}=0$.
5.1.2 Numerical performance of Perishing-Guardrail
Empirical Trade-off Curves.
@@ 5.1.2 Numerical performance of Perishing-Guardrail
We numerically investigate the impact of the perishing rate on the envy-efficiency frontier, and use the empirical trade-off curve to provide guidance on how decision-makers should select $L_{T}$ in the setting of geometric i.i.d. perishing. The demands $N_{t}$ are drawn from a truncated normal distribution $\mathcal{N}(2,0.25)$. We let $T=100$, $B=200$, and vary $\alpha\in\{0.1,0.2,0.25,0.3\}$ (see Appendix D, Fig. 7 for additional values of $\alpha$). For these instances, we compare Perishing-Guardrail and Vanilla-Guardrail (Sinclair et al., , 2022), a guardrail-based algorithm designed for settings without perishable resources, with $L_{T}=T^{-\beta}$, $\beta\in\{0,0.05,0.1,\ldots,1\}$. All results are averaged over 150 replications.
@@
The empirical trade-off curves can be found in Fig. 4. For $\alpha\in\{0.1,0.2\}$, Vanilla-Guardrail makes close to no gains in efficiency for any value of $L_{T}$. In these high-perishing settings, then, setting $L_{T}=0$ is optimal (there is no trade-off), in stark contrast to the classic setting without perishability. As $\alpha$ increases, Vanilla-Guardrail attains small gains in efficiency, but plateaus very quickly. This yields the important insight that perishing-agnostic algorithms are not able to leverage unfairness to improve efficiency in the presence of perishable resources. Perishing-Guardrail, on the other hand, sees extremely large gains in efficiency for a very small increase in envy, across all values of $\alpha$. Even for this algorithm, however, larger values of $L_{T}$ do not provide marginal gains in efficiency; the horizontal asymptote observed across all values of $\alpha$ is precisely the cumulative $\sigma$-induced loss, $T\mathcal{L}^{\textsf{perish}}$. Moreover, the vertical asymptote across all plots corresponds to the unavoidable loss due to demand uncertainty (Theorem 3.7).
@@
Note that, for small values of $\alpha$, the Perishing-Guardrail empirical trade-off curve lies to the left of that of Vanilla-Guardrail, i.e., it achieves lower counterfactual envy across all values of $L_{T}$. As we will see below, this is due to the fact that Vanilla-Guardrail achieves extremely high stockout rates. This effect is diminished as $\alpha$ increases (i.e., the perishing rate decreases). As this happens, both curves move down and to the left (and closer) as they achieve lower counterfactual envy and inefficiency due to spoilage. When the perishing rate is negligible (largest value of $\alpha$), the empirical trade-off curve of Vanilla-Guardrail is slightly to the left of that of Perishing-Guardrail; this is due to the loss incurred by our modified $\underline{X}$ construction, which always allocates less than $B/\overline{N}$ as a baseline. However, even when perishing is negligible Perishing-Guardrail is slightly more efficient that Vanilla-Guardrail, despite its baseline allocation being lower. This runs counter to the intuition that Vanilla-Guardrail should be more efficient since it has a higher baseline allocation. The reason for this is the difference in the two algorithms’ threshold allocation decisions. Our algorithm, Perishing-Guardrail, allocates $\overline{X}=\underline{X}+L_{T}$ if it forecasts that it has enough budget remaining to allocate $\overline{X}$ in period $t$, and $\underline{X}$ onwards. On the other hand, Vanilla-Guardrail allocates $B/\overline{N}+L_{T}$ if it has enough budget remaining to allocate this high amount in period $t$, and $B/\overline{N}$ in all future periods. Since $B/\overline{N}>\underline{X}$, Vanilla-Guardrail depletes its budget faster than Perishing-Guardrail whenever they both allocate the lower guardrail. Hence, Perishing-Guardrail is able to allocate aggressively more frequently than Vanilla-Guardrail, which results in improved efficiency and decreased spoilage.
@@
These empirical trade-off curves help to provide guidance on the choice of $L_{T}$. In particular, across all experiments, the cusp of the trade-off curve lies at $L_{T}\sim T^{-0.35}$. This value of $L_{T}$ is larger than no-perishing cusp of $L_{T}\sim T^{-1/2}$ (Sinclair et al., , 2022). This is due to the fact that our baseline allocation is significantly lower to avoid perishing-induced stockouts; hence, in order to recover from this inefficiency, $L_{T}$ must be higher. We use this observation in the following experiments, comparing the performance of Perishing-Guardrail and Vanilla-Guardrail for $L_{T}\sim T^{-0.35}$.
Perishing-Guardrail performance: synthetic experiments.
We compare the performance of Perishing-Guardrail to three benchmark algorithms:
  • Vanilla-Guardrail (Sinclair et al., (2022));
  • Static-$\underline{X}$, the algorithm which allocates $X_{t,\theta}=\underline{X}$ for all $t,\theta$, until it runs out of resources;
  • Static-$\frac{B}{\overline{N}}$, the algorithm which allocates $X_{t,\theta}=\frac{B}{\overline{N}}$ for all $t,\theta$, until it runs out of resources.
We are interested in five metrics: $(i)$ expected counterfactual envy $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it EF}}}\right]$, $(ii)$ expected hindsight envy $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Envy}}\right]$, $(iii)$ expected inefficiency $\mathbb{E}\mathopen{}\mathclose{{}\left[\Delta_{\text{\it efficiency}}}\right]$, $(iv)$ expected spoilage $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{PUA}_{\geq 1}}\right]$, and $(v)$ the stockout probability $\mathbb{E}\mathopen{}\mathclose{{}\left[\textsc{Stockout}}\right]$, i.e. the proportion of replications for which the algorithm runs out of resources before the end of the time horizon. We use the same simulation setup as above, with $B=2T$ (see Appendix D, Fig. 8 for additional values of $\alpha$). Fig. 5 illustrates the performance of each algorithm across the first four metrics of interest. We identify three regimes:

### v2

@@ 4.3 Analysis
Case 2: $\underline{X}\leq\frac{B}{N}\leq\overline{X}$. We have: $|X_{t,\theta}^{alg}-B/N|=\max\left\{\frac{B}{N}-\underline{X},\overline{X}-\frac{B}{N}\right\}\leq\overline{X}-\underline{X}=L_{T}.$
@@
where the inequality follows from $X_{t_{0}}^{alg}=\underline{X}$ and the threshold condition for allocating $\overline{X}$.
@@
$\leq 2\beta_{avg}\sqrt{2\tilde{c}|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}\min\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}\right\},$ (18)
@@
Finally, consider $\overline{P}_{t_{0}}+\textsc{PUA}_{<t_{0}}^{alg}$. By construction, $\overline{P}_{t_{0}}\leq\overline{P}_{1}=\overline{\Delta}(\underline{X})$. To upper bound $\textsc{PUA}_{<t_{0}}^{alg}$, we consider the process that allocates $\underline{X}$ in each period to all arrivals. Let $B_{t}(\underline{X})$ denote the quantity of remaining items under this process, and $\mathcal{B}_{t}(\underline{X})$ the set of remaining items. We use $\textsc{PUA}_{t}(\underline{X})$ to denote the quantity of unallocated items that perish at the end of period $t$ under this process, and $\textsc{PUA}_{<t}(\underline{X})$ those that perished before the end of period $t$. The following lemma allows us to tractably bound $\textsc{PUA}_{<t_{0}}^{alg}$ via this process. We defer its proof to Appendix 8.3.4.
Lemma 4.13
@@
where the second inequality follows from $\overline{\Delta}(\underline{X})\leq B-\overline{N}\underline{X}=B-\overline{N}\left(\frac{B}{\overline{N}}-\mathcal{L}^{\textsf{perish}}\right)=\overline{N}\mathcal{L}^{\textsf{perish}}$, and the last inequality uses the definition of $\overline{N}$. Putting bounds (4.12), (19) and (20) together, we obtain:
$\Delta_{\text{\it efficiency}}$ $\leq 2\beta_{avg}\sqrt{2\tilde{c}|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}\min\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}\right\}+L_{T}|\Theta|(\mu_{max}+\rho_{\max})$
@@
[…]
Remark 4.14
While Perishing-Guardrail achieves the optimal guarantees for arbitrary $\sigma$ and perishing distributions, the perishing-agnostic guardrail algorithm of Sinclair et al. (2023) is optimal in a special case of our setting. Namely, suppose the decision-maker knows the hindsight optimal ordering, and that there exist deterministic constants $\overline{N}$ and $(\overline{P}_{<t},\underline{N}_{<t},t\geq 2)$, such that
$\mathbb{P}\left(P_{<t}\leq\overline{P}_{<t},N_{<t}\geq\underline{N}_{<t},N\leq\overline{N}\ \forall\ t\geq 2\right)\geq 1-\delta\quad\text{and}\quad\frac{\overline{P}_{<t}}{B}\leq\frac{\underline{N}_{<t}}{\overline{N}}\quad\forall\ t\ \geq 2,$ (21)
for $\delta\in(0,1)$. Under the above condition, offset-expiry holds with probability at least $1-\delta$. Hence, $\mathcal{L}^{\textsf{perish}}=0$ over the sample paths for which (21) holds. As a result, the baseline allocation $\underline{X}=\frac{B}{\overline{N}}$ is feasible over these sample paths. Since allocating aggressively relative to $\underline{X}$ cannot increase the amount of goods that spoil, it follows that the algorithm of Sinclair et al. (2023) is optimal in this setting. While the assumption that the hindsight optimal ordering is known is unlikely to hold in practical settings where the perishing times are stochastic, this special case further highlights the two main features that render this setting challenging: $(i)$ the inherent aggressiveness of the perishing process, and $(ii)$ the unknown perishing order.
4.4 Special Cases
We conclude the section by instantiating our bounds for important special cases. Recall, Theorem 3.3 established that the proportional allocation $B/N$ is achievable if and only if the process is offset-expiring. As a result, for general, non-offset-expiring perishing processes, one should expect $\mathcal{L}^{\textsf{perish}}=\Omega(1)$. In this case, Corollaries 4.2 and 4.3 imply that counterfactual envy is constant, and inefficiency grows linearly in $T$. In this section we restrict our attention to a special class of distributions, parameterized by $\delta\in(0,1)$, which we term $\delta$-offset-expiring, for which $\mathcal{L}^{\textsf{perish}}$ vanishes with respect to $T$ under the hindsight optimal ordering. We formally define this class of processes below.
Definition 4.15 ($\delta$-offset-expiry.)
Fix $\delta\in(0,1)$. A process is $\delta$-offset-expiring if:
${\mathbb{P}\left(\frac{P_{<t}}{B}\leq\frac{N_{<t}}{N}\ \forall\,t\geq 2\right)\geq 1-\delta.}$ (22)
For simplicity, in this section we assume $B=N=T$, with $N_{t}=1$ for all $t$, almost surely. [^2] At the cost of cumbersome algebra, one can relax this assumption and derive entirely analogous results. In this case, Equation 22 reduces to the condition that $\mathbb{P}(P_{<t}\leq t-1\ \forall\ t\geq 2)\geq 1-\delta$.
4.4.1 The i.i.d. setting.
We introduce some additional notation. For $t\in[T]$, we let $\mathcal{B}^{rand}_{<t}=\{b:\mathbb{P}\left(T_{b}<t\right)\in(0,1)\}$ and $\mathcal{B}^{det}_{<t}=\{b:\mathbb{P}\left(T_{b}<t\right)=1\}$. In words, $\mathcal{B}^{det}_{<t}$ represents the set of items that perish before period $t$, almost surely.
Propositions 4.16 and 4.17 respectively establish necessary and sufficient conditions for $\delta$-offset-expiry to hold when the perishing distribution is i.i.d., for non-trivial values of $\delta$.
Proposition 4.16
Suppose there exists $t\geq 2$ such that $\mathbb{E}[P_{<t}]>t-1$. If $\mathcal{B}^{rand}_{<t}=\emptyset$, $\delta$-offset-expiry cannot be satisfied for any value of $\delta\in(0,1)$. Else, $\delta$-offset-expiry cannot be satisfied for $\delta<\frac{1}{2}-{\mathrm{Std}\left[P_{<t}\right]^{-3}}\cdot T$.
Proposition 4.17
Suppose $\mathbb{E}[P_{<t}]\leq t-1$ for all $t\geq 2$. Then, the perishing process is $\delta$-offset-expiring for any $\delta\geq\sum_{t=2}^{T}\min\left\{\left(\frac{\mathrm{Std}\left[P_{<t}\right]}{t-\mathbb{E}[P_{<t}]}\right)^{2},\exp\left(-\frac{{2}(t-\mathbb{E}[P_{<t}])^{2}}{|\mathcal{B}^{rand}_{<t}|}\right)\right\}\mathds{1}\{|\mathcal{B}^{rand}_{<t}|>0\}.$
Propositions 4.16 and 4.17 together illustrate how the mean and variability of the perishing distribution jointly impact offset-expiry. In particular, Proposition 4.17 states that either the expected lag between demand and perishing, given by $t-\mathbb{E}[P_{<t}]$, must be large, or the coefficient of variation (CV) with respect to the random lag process $t-P_{<t}$, must be small. This then reduces to a bound on the variance of $P_{<t}$ in settings where perishing closely tracks demand.
Additionally, note that the necessary condition in Proposition 4.16 fails to hold for one of the most standard models of perishing: geometrically distributed perishing with parameter $1/T$ (that is, a constant fraction of items perish in each period). This highlights that one of the most popular models in the literature is, in a sense, far too pessimistic; for this setting, there is no hope of achieving low envy and efficiency with high probability. Proposition 4.18 below shows how the perishing rate under this popular distribution affects offset-expiry and the perishing-induced loss $\mathcal{L}^{\textsf{perish}}$.
Proposition 4.18
Suppose $T_{b}\sim\emph{Geometric}(p)$ for all $b\in\mathcal{B}$, with $p=o(1/T)$. Then, the perishing process is $\delta$-offset-expiring for any $\delta\geq 2\log T\cdot\frac{Tp}{(1-Tp)^{2}}$. Moreover, $\underline{X}\geq 1-3Tp-\frac{\log(3\log(T)/\delta)}{T}$ for any ordering $\sigma$. As a result,
$\mathcal{L}^{\textsf{perish}}\leq 3Tp+\frac{\log(3\log(T)/\delta)}{T}.$
Figure 4: Maximum feasible allocation $\underline{X}$ vs. $\alpha$, for $T_{b}\sim\emph{Geometric}(T^{-(1+\alpha)})$, $B=300$, $T=150$, $N_{t}$ drawn from a truncated normal distribution $\mathcal{N}(2,0.25)$, and $\delta=1/T$. Here, $\underline{X}$ was calculated via line search, with Monte Carlo simulation used to estimate $\Delta(X)$ for each value of $X$. The dashed line represents the “naive” allocation $B/\overline{N}=0.89$ which ignores possible perishing, and the green line is the curve of best fit to $\underline{X}$.
Proposition 4.18 establishes that, in the worst case, $\underline{X}$ decays linearly in the rate $p$ at which goods spoil. This highlights the extent to which the commonly used (and practically pessimistic) geometric model limits both the kinds of perishable goods selected before allocation, as well as the rate at which a decision-maker can allocate goods. Letting $p=T^{-(1+\alpha)}$, $\alpha\in(0,1)$, Proposition 4.18 implies that $\mathcal{L}^{\textsf{perish}}$ is on the order of $T^{-\alpha}$. Alternatively, if a decision-maker wants no more than $T^{-\alpha}$ loss relative to the proportional allocation, Proposition 4.18 provides an upper bound of $T^{-(1+\alpha)}$ on $p$.
We illustrate the scaling of $\underline{X}$ under geometric perishing numerically in Figure 4, for an instance where the decision-maker also faces demand uncertainty. We observe that $\underline{X}$ is concave increasing in $\alpha$, and that the lower bound in Proposition 4.18 provides a good fit for $\underline{X}$, even under demand uncertainty. For $\alpha$ close to 0 (i.e., $p\sim 1/T$), $\underline{X}\approx 0.4$, less than half of the “naive” no-perishing allocation, $B/\overline{N}=0.89$. For $\alpha=1$, $\underline{X}\approx B/\overline{N}$. Note that, for $\alpha>1$, $\underline{X}$ is limited by the confidence bound $\log(3\log T/\delta)/T$ in Proposition 4.18. Plugging the lower bound on $\delta$ into this term, this implies that, even under deterministic demand, $\mathcal{L}^{\textsf{perish}}=\Omega(1/T)$.
4.4.2 The non-i.i.d. setting.
When the perishing distribution is not identical across different units, the quality of the allocation schedule $\sigma$ can have a significant impact on the perishing-induced loss $\mathcal{L}^{\textsf{perish}}$. Proposition 4.19 illustrates how the allocation schedule and perishing distribution come together in determining $\mathcal{L}^{\textsf{perish}}$.
Proposition 4.19
Suppose there exists $\alpha\in(0,1)$ such that:
  1. 1.
    $\mathbb{E}\left[T_{b}\right]>\min\left\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\right\}$
  2. 2.
    $\sum_{b}\frac{\mathrm{Var}\left[T_{b}\right]}{\left(\mathbb{E}\left[T_{b}\right]-\min\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\}\right)^{2}}\leq\frac{1}{2}T^{1-\alpha}$.
Then, for any $\delta\geq 3\log(T)e^{-\frac{1}{8}T^{1-\alpha}}$, the process is $\delta$-offset-expiring, and $\underline{X}\geq 1-T^{-\alpha}$. As a result, $\mathcal{L}^{\textsf{perish}}\leq T^{-\alpha}$.
Proposition 4.19 highlights the two key components that determine the baseline allocation: $(i)$ the lag between the expected perishing time $\mathbb{E}\left[T_{b}\right]$ and the expected allocation time $\min\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\}$ (which we colloquially refer to as “room to breathe”), and $(ii)$ the variance of the perishing time. Specifically, Condition 1 implies that, if $\mathbb{E}\left[T_{b}\right]$ is low, it must be that the item is allocated early on in the horizon (i.e., $\sigma(b)$ is low). This encodes the “race against time” intuition that is typically held around perishing. Condition 2 can be viewed as an upper bound on the cumulative adjusted coefficient of variation of the perishing process. High-variance perishing times and smaller “room to breathe” push $\alpha$ down, resulting in a lower maximum feasible allocation rate, and consequently higher perishing-induced loss. In Section 5.4 we build on this intuition to propose heuristic schedules with low $\mathcal{L}^{\textsf{perish}}$.
@@ 5 Numerical Experiments
In this section we study the practical performance of Perishing-Guardrail via an extensive set of numerical experiments. We first consider one of the most popular (and aggressive) models of perishing: geometrically distributed perishing times. We explore the dependence of the envy-efficiency trade-off on the perishing rate for this setting, leveraging the empirical trade-off curves to provide guidance on the choice of envy parameter $L_{T}$. We moreover compare the performance of Perishing-Guardrail to its perishing-agnostic counterpart (Sinclair et al. 2023) on both synthetic and real-world datasets (Keskin et al. 2022). Lastly, for the non-i.i.d. setting we numerically study choices of allocation schedules that achieve the two-fold desideratum of being practical for decision-makers and achieving small perishing-induced loss. See https://github.com/seanrsinclair/perishing_fairness for the code database.
5.1 Description of Metrics and Benchmark Policies
We evaluate Perishing-Guardrail using $\delta=\frac{1}{T}$, benchmarking it against three algorithms from the literature:
  • Vanilla-Guardrail (Sinclair et al. (2023)): a guardrail-based algorithm designed for settings without perishable resources;
  • Static-$\underline{X}$: the algorithm that allocates $X_{t,\theta}=\underline{X}$ for all $t,\theta$, until it runs out of resources;
  • Static-$\frac{B}{\overline{N}}$: the algorithm that allocates $X_{t,\theta}=\frac{B}{\overline{N}}$ for all $t,\theta$, until it runs out of resources.
We report the performance of these algorithms along five metrics: $(i)$ expected counterfactual envy $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$, $(ii)$ expected hindsight envy $\mathbb{E}\left[\textsc{Envy}\right]$, $(iii)$ expected inefficiency $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$, $(iv)$ expected spoilage $\mathbb{E}\left[\textsc{Spoilage}\right]$, and lastly $(v)$ stockout probability $\mathbb{E}\left[\textsc{Stockout}\right]$, i.e., the proportion of replications for which the algorithm runs out of resources before the end of the time horizon.
5.2 Geometric Perishing: Synthetic Experiments
@@
We first consider the setting in which each unit perishes independently with probability $p$ in each period, i.e., $T_{b}\sim\text{Geometric}(p)$, for all $b\in\mathcal{B}$. Throughout the section, we assume $|\Theta|=1$. Since perishing times are identically distributed, the allocation order $\sigma$ does not have any impact on the performance of the algorithm; hence, we assume $\sigma$ is the identity ordering. To study the impact of the perishing rate on algorithm performance, we let $p=T^{-(1+\alpha)}$, varying $\alpha\in\{0.1,0.2,0.25,0.3\}$ (see Section 9 for additional values of $\alpha$). We let $T=100$, $B=2T$, and draw demand $N_{t}$ from a truncated normal distribution $\mathcal{N}(2,0.25)$, for $t\in[T]$. All results are averaged over 150 replications.
Empirical trade-off curves.
@@
Figure 5: Empirical trade-off between $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$ and $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$. The points on the trade-off curve correspond to increasing values of $L_{T}$, from left to right. Static-$\frac{B}{\overline{N}}$ and Static-$\underline{X}$ respectively correspond to Vanilla-Guardrail and Perishing-Guardrail for $L_{T}=0$.
@@
We begin by numerically investigating the impact of the perishing rate on the envy-efficiency frontier, and use the empirical trade-off curves to provide guidance on how decision-makers should select $L_{T}$. For these instances, we compare Perishing-Guardrail and Vanilla-Guardrail (Sinclair et al. 2023) with $L_{T}=T^{-\beta}$, $\beta\in\{0,0.05,0.1,\ldots,1\}$.
@@
The empirical trade-off curves for each value of $\alpha$ can be found in Figure 5. In each case, we report the probability that a sample path is offset-expiring, denoted by $\mathbb{P}(\mathcal{E}_{OE})$. For $\alpha\in\{0.1,0.2\}$ (Figures 5(a) and 5(b)), Vanilla-Guardrail makes close to no gains in efficiency for any value of $L_{T}$. In these high-perishing settings, then, setting $L_{T}=0$ is optimal (there is no trade-off), in stark contrast to the classic setting without perishability. As $\alpha$ increases (Figures 5(c) and 5(d)), Vanilla-Guardrail attains small gains in efficiency, but plateaus very quickly. This yields the important insight that perishing-agnostic algorithms are not able to leverage unfairness to improve efficiency in the presence of perishable resources. Perishing-Guardrail, on the other hand, sees extremely large gains in efficiency for a very small increase in envy, across all values of $\alpha$. For larger values of $L_{T}$, we observe marginal gains in efficiency; the horizontal asymptote observed across all values of $\alpha$ is precisely the cumulative perishing-induced loss, $T\mathcal{L}^{\textsf{perish}}$ (Theorem 3.8). Moreover, the vertical asymptote across all plots corresponds to the unavoidable loss due to demand uncertainty (Theorem 2.4).
@@
Note that, for small values of $\alpha$, the Perishing-Guardrail empirical trade-off curve lies to the left of that of Vanilla-Guardrail, i.e., it achieves lower counterfactual envy across all values of $L_{T}$. As we will see below, this is due to the fact that Vanilla-Guardrail achieves extremely high stockout rates. This effect is diminished as $\alpha$ increases (i.e., the perishing rate decreases). As this happens, both curves move down and to the left (and closer) as they achieve lower counterfactual envy and inefficiency due to spoilage. When the perishing rate is negligible ($\alpha=0.3$), the empirical trade-off curve of Vanilla-Guardrail is slightly to the left of that of Perishing-Guardrail; this is due to the loss incurred by our modified $\underline{X}$ construction, which always allocates less than $B/\overline{N}$ as a baseline. However, even when perishing is negligible Perishing-Guardrail is slightly more efficient that Vanilla-Guardrail, despite its baseline allocation being lower. This runs counter to the intuition that Vanilla-Guardrail should be more efficient since it has a higher baseline allocation. The reason for this is the difference in the two algorithms’ threshold allocation decisions. Our algorithm, Perishing-Guardrail, allocates $\overline{X}=\underline{X}+L_{T}$ if it forecasts that it has enough budget remaining to allocate $\overline{X}$ in period $t$, and $\underline{X}$ onwards. On the other hand, Vanilla-Guardrail allocates $B/\overline{N}+L_{T}$ if it has enough budget remaining to allocate this high amount in period $t$, and $B/\overline{N}$ in all future periods. Since $B/\overline{N}>\underline{X}$, Vanilla-Guardrail depletes its budget faster than Perishing-Guardrail whenever they both allocate the lower guardrail. Hence, Perishing-Guardrail is able to allocate aggressively more frequently than Vanilla-Guardrail, which results in improved efficiency and decreased spoilage.
@@
These empirical trade-off curves moreover help to provide guidance on the choice of $L_{T}$. In particular, across all experiments, the cusp of the trade-off curve lies at $L_{T}\sim T^{-0.35}$, which is larger than no-perishing cusp of $L_{T}\sim T^{-1/2}$ (Sinclair et al. 2023). This is due to the fact that our baseline allocation is significantly lower to avoid perishing-induced stockouts; hence, in order to recover from this inefficiency, $L_{T}$ must be higher. We use this observation in the following experiments, comparing the performance of Perishing-Guardrail and Vanilla-Guardrail for $L_{T}\sim T^{-0.35}$.
Benchmark comparisons.
Next, we compare the performance of all algorithms along our five metrics as the number of rounds $T$ grows large, for $\alpha\in\{0.1,0.2,0.3\}$. Based on our prior insights, we use $L_{T}=T^{-0.35}$ for both Perishing-Guardrail and Vanilla-Guardrail. Our results can be found in Figure 6.